OpenStudy (anonymous):

Integrate sin(2t+pi)

OpenStudy (tkhunny):

Integrate sin(2t).

jimthompson5910 (jim_thompson5910):

hint: let u = 2t+pi derive both sides with respect to t to get du/dt = 2. That turns into du = 2*dt and dt = du/2

OpenStudy (tkhunny):

...or, now that I think about it... \(\sin(2t+\pi) = \sin(2t)\cos(\pi) + \cos(2t)\sin(\pi) = -\sin(2t) = -2\sin(t)\cos(t)\) Pretty obvious substitution going on, there.